I just want to know if I am solving the problems right.
1) Solve each of the following for the requested variable:
a) P=2L+2W for L
P-2W=2L+2W-2W
P-2W=2L/2
P-2W/2=L

b)(3x+Y)/2=2 for y
3x+y=2*2
3x+y=4
3x-3x+y=4-3x
y=4-3x
2)The linear equation 3x + 4y = 12 is in Slope intercept form.
3)What is the slope of 3x+4y=12. The slope is 3.

Answer by Alan3354(69443)

I just want to know if I am solving the problems right.
1) Solve each of the following for the requested variable:
a) P=2L+2W for L
P-2W=2L+2W-2W
P-2W=2L/2
P-2W/2=L
-------------
p = 2L + 2W
2L = P - 2W
L = (P-2W)/2
-------------
b)(3x+Y)/2=2 for y
3x+y=2*2
3x+y=4
3x-3x+y=4-3x
y=4-3x Looks good.
-----------
2)The linear equation 3x + 4y = 12 is in Slope intercept form.
It's not, y = mx + b is slope-intercept form.
---------------------
3)What is the slope of 3x+4y=12. The slope is 3.
Put into slope-intercept form (that means solve for y)
4y = -3x + 12
y = (-3/4)x + 3
Slope = -3/4